Christmas Express makes wreaths in batch sizes of 30. The cutting & assembly process takes 4 minutes per wreath and the decorating process time is 15 minutes per wreath. It takes 4 minutes to move the wreaths from the cutting & assembly process to the decorating process.

Compute the value-added, non-value-added, and the total lead time of the wreath process.

Value-added lead time:
Non-value-added lead time:
Move time
Within batch wait time
Total lead time

Compute the value-added ratio. Round to one decimal place.
%
Answer #1

(1)

Value added lead time = cutting & assembly process time + decorating process time

= 4 + 15 = 19 minutes

(2)

Non-value added lead time:

Move time = 4 minutes

within batch wait time = (total time to perform operations) X ( batch size - 1)

= ( 4 + 15) X (30 - 1) = 551 minutes

(3)

Total lead time = 19 + 4 + 551 = 574 minutes

(4)

Value added ratio = Value added lead time / Total lead time Value added ratio

= 19/574 = 3.3%